The following record came from the University Registrar at O.A.U, Ile-Ife some years ago. The record showed that the total enrolment in the university was 10,000 students. Of these, the Registrar stated that 2,521 were married, 6,471 were men, 3,115 were over 21 years of age, 1,915 were married men, 1,873 were married persons over 21 years and 1,302 were married men over 21 years of age. 512 women were over 21 years and not married while 3,826 men were not married and not over 21 years.
Find the difference between the women that are not married and not over 21 years and the men that are not married but over 21 years.

Answer by Theo(13342)

10000 total students.
6471 were men which means 3529 were women.

2521 were married which means that 10000 - 2521 = 7479 were not married.

3115 were over the age of 21 which means that 10000 - 3115 = 6885 that were not over the age of 21.

1915 were married men which means that 2521 - 1915 = 606 were married women which means that 3529 - 606 = 2923 were women who were not married.

1915 were married men which means that 6471 - 1915 = 4556 were men who were not married.

3826 men were not married and not over 21 which means that 4556 - 3826 = 730 were men who were not married and over 21.

1873 were married persons over 21 which means that 2521 - 1873 = 648 were married persons not over 21.

1302 were married men over 21 which means that 1915 - 1302 = 613 were married men not over 21.

512 women were over 21 and not married which means that 2923 - 512 = 2411 were women who were not married and not over 21.

the number of women who were not married and not over 21 = 2411.
the number of men who were not married and over 21 = 730.

the difference between the women that are not married and not over 21 years and the men that are not married but over 21 years = 2411 - 730 = 1681.

the following chart shows the relationship.
from this chart, you should be able to derive all the givens.

those givens are:

10000 total students.
2521 married.
6471 men
3115 over the age of 21
1915 married men
1873 married persons over 21
1302 married men over the age of 21
512 women over 21 and not married
3826 men not over 21 and not married

$$$